• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Templon

    Log entry time 08:07:20 on October 3,2000

    Entry number 49165

    keyword=graveyard shift summary

    After 4He elastics, stopped so Mark Stevens could change polarity
    of spectrometer on right side.
    run 1344 is unknown state; not sure if this run was started
    before or after we turned beam off and started to drop SR momentum.
    polarity change takes longer than expected since JAT did not
    think about implications of changing SR momentum (takes a long
    time). Mark came in and had to sit around for 2 hrs waiting for
    the field to die.
    Mark Stevens leaves at 4:10. Cycle up quads Q2 & Q3 in SR to 1400 A.
    also cycling quads in SL to 1600 A (mom must increase here too)
    Angle of SR set to 19.853 deg (19.5 floor, 28.5 vernier). p0 of spectrometers
    are still setting (hadron takes a long time). sawatzky is downloading
    the proper coincidence trigger.
    5:15 - beam on target. SR field still coming up but we want to look
    .. for coincidences. also realized that we may need to download new HV
    .. settings for SR now that polarity is changed. starting run 1345 to look
    .. at spot size. spot looks great.
    i just realized that we will not see any coincs here since we are set up for
    ... delta=0 for elastics in the electron arm. furthermore, it is mandatory
    ... that we have the proper HV settings here for SR, since this is a normalization
    ... run. so we will look for the proper HV settings.
    cannot find them. will assume that the settings we have are OK.
    SR dipole is crawling. we set the momentum for SR at 4:25 and it
    looks like (at 5:50) it will take another half hour to reach the target
    (if it does not continue to increase more slowly as time passes, which
    has been the case ... if this keeps up we will get there tomorrow).
    we are going to set the field a bit above, and then come back down
    to the desired setting.
    www.jlab.org is down so we cannot access HALOG. so we cannot look
    for old HV settings for the hadron arm. also since some PMTs were
    replaced it is not clear that this is a good thing to use anyway.
    run 1346 is a run with pedestal suppression turned back on (restoreped.com)
    SR field still (6:15) changing. ask for 2 uA.
    brad saw that in SL, NMR regulation was ON, but in SR NMR regulation was
    OFF. We turned it ON and now the search for p0 seems to be progressing
    in a normal fashion.
    starting first link point -- 4He elastics SL, PY-1A setting SR.
    run 1347 - trial run at 20 uA. good run but not enough statistics.
    run 1348 - trial run at 40 uA. stop to increase prescale factors.
    run 1348 - run at 40 uA. prescale factors again
    run 1349 - run at 40 uA. good run
    run 1350 - beam unstable, tripped off. not really a good run.